OpenStudy (anonymous):

ILL MEDAL/FAN (NO DIRECT ANSWERS, PLZ) :) When the function f(x) = 3(5)^x is changed to f(x) = 3(5)^x + 22, what is the effect? There will be no change to the graph because the exponential portion of the function remains the same. The y-intercept is 22 spaces higher. The x-intercept is 22 spaces higher. All input values are moved 22 spaces to the right

OpenStudy (anonymous):

@DanJS @Hero @Afrodiddle

OpenStudy (anonymous):

@Michele_Laino @Miracrown

OpenStudy (anonymous):

I think its C :)

OpenStudy (anonymous):

Maybe B? I am just not sure :(

OpenStudy (anonymous):

@ganeshie8

OpenStudy (jdoe0001):

\(\bf \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards\\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ % template start \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ 3}} (5)^{( x + {\color{red}{ C}} )} + {\color{green}{ D}}\\ %template end \qquad\qquad \quad \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ {\color{red}{ C}}\end{array} \begin{array}{llll}{\color{red}{ C}} > 0 & to\ the\ left\\ {\color{red}{ C}} < 0& to\ the\ right\end{array} \end{array}\) what do you think?

OpenStudy (anonymous):

I think that tthe answer is B

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

is it? what makes you think is B?

OpenStudy (anonymous):

because f(x) is total y, and nothing else is changing.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Its either B or C, I know that much. @jdoe0001

OpenStudy (anonymous):

sorry my connection lost :( but I'm back! :D

OpenStudy (anonymous):

lol all good :)

OpenStudy (jdoe0001):

so.. ok... why is it B or C though?

OpenStudy (anonymous):

Hmmm I'm not quite sure how to explain it to you, but is x the exponent?

OpenStudy (anonymous):

because it says +22... arrgghh

OpenStudy (anonymous):

and yes lol

OpenStudy (anonymous):

lol my "bedtime" is soon, so B sounds right? Thats what I thought..

OpenStudy (jdoe0001):

ok now, notice the function transformations \(\bf \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards\\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ % template start \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ 3}} (5)^{( x + {\color{red}{ C}} )} + {\color{green}{ 22}}\\ %template end \qquad\qquad \quad \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ {\color{red}{ C}}\end{array} \begin{array}{llll}{\color{red}{ C}} > 0 & to\ the\ left\\ {\color{red}{ C}} < 0& to\ the\ right\end{array} \end{array}\) notice the vertical shift

OpenStudy (anonymous):

It cant be c, since the x value itself is not affected by the 22.

OpenStudy (anonymous):

so its BBBBBBBBBBBB!

OpenStudy (anonymous):

@jdoe0001 @mathway

OpenStudy (jdoe0001):

correct, the +22 in the function transformation, refers to a vertical shift, D > 0, means is upwards so 'y" is up bu 22 units

OpenStudy (jdoe0001):

up rather, by 22 units

OpenStudy (anonymous):

thankkkksss!! :) Good night!

OpenStudy (jdoe0001):

yw